Average daily demand (d) 130 cases
Demand Standard deviation (σd) 20 cases
Service level 94.0%
Service level alternative 86.0%
Lead tme, LT 6 days

1. What is the standard deviation of demand during lead time? Round your answer to one decimal place.

_______ cases

2. What is the required safety stock at a 94 percent service level? Round your answer to the nearest whole number.

_______ cases

3. What is the reorder point? Round your answer to the nearest whole number.

_______ cases

4.What is the required safety stock at a 86 percent service level? Round your answer to the nearest whole number.

_______ cases

5. What is the safety stock reduction? Round your answer to the nearest whole number.

_______ cases
Answer #1

SD of demand during lead time = SD*(LT)^0.5

= 20*(6)^0.5= 49

2. Safety stock = z* SD during LT

= 1.55*49= 76

3.ROP = demand during lead time +safety stock

= 130*6+76 = 856

4. SS at 86% SL.= 1.08*49= 53

Difference in SS = 76-53=23